A box contains 5 green and 2 blue dice.Four dice are chosen one after the other .Determine the values of the random variable G representing the number of green dice.


1
Expert's answer
2022-02-14T17:46:51-0500

The possible values of the random variable "G" representing the number of green dice are

"2, 3, 4."


"P(G=2)=\dfrac{\dbinom{5}{2}\dbinom{2}{4-2}}{\dbinom{5+2}{4}}=\dfrac{10(1)}{35}=\dfrac{2}{7}""P(G=3)=\dfrac{\dbinom{5}{3}\dbinom{2}{4-3}}{\dbinom{5+2}{4}}=\dfrac{10(2)}{35}=\dfrac{4}{7}""P(G=4)=\dfrac{\dbinom{5}{4}\dbinom{2}{4-4}}{\dbinom{5+2}{4}}=\dfrac{5(1)}{35}=\dfrac{1}{7}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 g & 2 & 3 & 4 \\ \hline
 P(g) & 2/7 & 4/7 & 1/7 \\
\end{array}"